OpenStudy (goformit100):

The magnitudes of two vectors A and B are A  5 units and B  2 units. Find the largest and smallest values possible for the resultant vector R =  A + B.

OpenStudy (nubeer):

maximum can be = 5+2 minimum can be = 5-2

OpenStudy (goformit100):

Thanks But How ?

OpenStudy (anonymous):

The maximum resultant is when both vectors are in the same direction and vice versa

Parth (parthkohli):



OpenStudy (anonymous):

In maximum case ,Just add them

Parth (parthkohli):



OpenStudy (goformit100):

Thankyou everyone .

OpenStudy (unklerhaukus):

the resultant will be a maximum when the vectors are parallel, and a minimum when the vectors are antiparallel